A runner wants to run 10,000m in 30 minutes. After 27 minutes he has run 8,900m. For how long must he accelerates at 0.02 m/s2 and then maintain that final speed?


1
Expert's answer
2021-02-22T10:20:52-0500
"t_{max}=(30-27)60=180\ s\\v=\frac{8900}{27\cdot60}=5.49\frac{m}{s}"

"10000=8900+5.49t+0.5(0.02)t^2\\t=156\ s"